OpenStudy (anonymous):

(1/12x^4+8/9x^3+5/8x^2+6)+(-7/12x^4+3/8x^2-6).... add

OpenStudy (anonymous):

http://www.youtube.com/watch?v=g5Lcb9KoIvI

OpenStudy (anonymous):

is this add? as in "combine like terms'?

OpenStudy (anonymous):

\[\frac{1}{12}x^4-\frac{7}{12}x^4=\frac{-6}{12}x^4=-\frac{1}{2}x^4\]

OpenStudy (anonymous):

\[\frac{8}{9}x^3\] has not match so we leave it be

OpenStudy (anonymous):

add

OpenStudy (anonymous):

\[\frac{5}{8}x^2+\frac{3}{8}x^2=\frac{8}{8}x^2=x^2\]

OpenStudy (anonymous):

apparently it wants the like terms in descending order

OpenStudy (anonymous):

\[6-6=0\]

OpenStudy (anonymous):

didn't know that until I tried to input my answer

OpenStudy (anonymous):

ok i just listed them for you. you just need to put it together

OpenStudy (anonymous):

\[-\frac{1}{2}x^4+\frac{8}{9}x^3+x^2\] i think that is all

OpenStudy (anonymous):

i don't get the descending order though :?

OpenStudy (anonymous):

btw it was king sighta, not king sighter

OpenStudy (anonymous):

descending order is how i put it. highest to lowest power

OpenStudy (anonymous):

i know that much but these things confuse me

OpenStudy (anonymous):

i only have one more shot to answer this one or it's going to give me a new problem

OpenStudy (anonymous):

by the way thanks for the youtube link